OpenStudy (nicnobles):

An equation of a circle is given as (x + 6)2 + (y − 7)2 = 81. Find the center and radius of the circle.

satellite73 (satellite73):

you don't "find" it, you "write it" \[(x-h)^2+(y-k)^2=r^2\] has center \((h,k)\) and radius \(r\)

OpenStudy (nicnobles):

@satellite73 thank you! could you help me with another?

satellite73 (satellite73):

sure

satellite73 (satellite73):

did you get the answer to that one?

OpenStudy (nicnobles):

@satellite73 yes i got (x + 6)^2 + (y – 7)^2 = 81, the center would be (-6, 7) and then the radius would be 9. is that correct?

satellite73 (satellite73):

yes

OpenStudy (nicnobles):

@satellite73 okay, thank you!

satellite73 (satellite73):

yw got another?

OpenStudy (nicnobles):

@satellite73 no i figured it out but thank you anyways!

satellite73 (satellite73):

ok good